maddybrown15:

A plane flying at a certain altitude is observed from two points that are 2.2 meters apart. The angles of elevation made by the two points are 55º and 72º. If the plane is to the east of both observation posts, the altitude of the plane is

Bearclaws72:

tan(55°) = x / (2.2 + y) tan (72°) = x / y You solve y from the second equation y = x / tan(72°) Substitute y in the first equation tan (55°) = x /(2.2 + x/tan(72°) Solve for x: x = tan(55)[2.2 + x/tan(72) ] x = tan(55) (2.2) + x tan(55)/tan(72) x - x tan (55)/tan(72) = 2.2 tan (55) Now substitute the values of the tangents tan (55) = 1.4281 tan(72) = 3.0777 x - 0.464x = 3.1419 x = 3.1419 / (1-0.464) =5.86 m